OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4) AB = ? sqrt ?

OpenStudy (anonymous):

sqrt 80

OpenStudy (anonymous):

whats the answer in simplest form

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[\sqrt{(-3-5)^2 +(4-8)^2}\] =\[4\sqrt{5}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

glad to help :)